OpenStudy (unknownrandom):

Find an equation of the tangent line to the curve y=2sinx+sin2x at the point (0,0).

OpenStudy (unknownrandom):

I can't figure this one out. Can someone talk me through it?

OpenStudy (solomonzelman):

You need to find the derivative first

OpenStudy (unknownrandom):

d/dx = -2 sin(x) - 2 sin(x) cos(x) sin(x)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=2\sin x+\sin 2x }\) \(\large\color{black}{ \displaystyle f'(x)=2\cos x+2\cos 2x }\)

OpenStudy (solomonzelman):

do you see how I got this?

OpenStudy (unknownrandom):

Yeah. For some reason I was using a different equation.

OpenStudy (solomonzelman):

d/dx [ 2sin(x) ] = 2 cos(x) d/dx [ sin(2x) ] = cos(2x) × {d/dx 2x] = cos(2x) × 2 = 2cos(2x)

OpenStudy (solomonzelman):

The derivative of the function is the function's slope. And thus, the derivative evaluated at x=0 (i.e f'(0)) is the slope of the function at x=0.

OpenStudy (unknownrandom):

The 2 in the second part is supposed to be to the second power.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=2\sin x+\sin^2x }\) this way?

OpenStudy (unknownrandom):

Yes

OpenStudy (unknownrandom):

\[2\cos(x)+2\sin(x)\cos(x)\]

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=2\sin x+(\sin x)^2 }\) \(\large\color{black}{ \displaystyle f'(x)=2\cos x+2(\sin x)^{2-1} \times \cos x}\)

OpenStudy (solomonzelman):

see what I did?

OpenStudy (unknownrandom):

Yep. So, after that you plug in 0 for x which makes the slope 2. Correct?

OpenStudy (jhannybean):

THat's right

OpenStudy (solomonzelman):

So you got: slope = 2 point: (0,0) do your line now:)

OpenStudy (unknownrandom):

Awesome. I was just using the wrong equation. Thanks guys!

OpenStudy (solomonzelman):

Anytime:)